OpenStudy (anonymous):

Find the real and complex zeros of the following function. Please show all of your work. f(x)=x^3-x^2+4x-4

OpenStudy (anonymous):

a good place to start is by trying x = 1, which gives you the sum of the coefficients

OpenStudy (anonymous):

\[1-1+4-4=0\] so 1 is a solution. therefore you can factor out the 1 to get \[x^3-x^2+4x-4=(x-1)(\text{something})\] and you can find the 'something" by either dividing or thinking

OpenStudy (anonymous):

diah40, look up "completing the square" on google. Once you re-remember (as we all have to do) how to do this, then finding the zeros will be a total cinch. For example, x(x-2)(x-3): zeros are: 0, 2, and 3.

OpenStudy (anonymous):

in any case it has to be \[x^3-x^2+4x-4=(x-1)(x^2+4)\] so only real zero is at x = 1, complex zeros at \[x=2i\] and its conjugate \[x=-2i\]

OpenStudy (anonymous):

Sorry, not completing the square, but factoring a cubic polynomial (or 3rd degree poly) http://www.youtube.com/watch?v=mXdAiJY7YfY